if x= 4 and y= -2, the value 0f 1/2xy^2 is? (^2 means squared).
Is the xy squared or just the y and how do you solve it?
(this problem is Question 17 on the Math A, August 2006 Packet @regentsprep.org).

Answer by checkley75(3666)

JUST THE Y IS SQUARED (1/2)*X*Y^2
(1/2)XY^2=(1/2)*4*(-2)^2=(1/2)*4*4=(1/2)*16=8